# Structured Packing in LLM Training Improves Long Context Utilization

### v8

@@ 3.3 Detailed study with medium models
In Table 5 and Table 5, we present a comprehensive examination of the impact of document packing on long-context performance using $270$M parameter models. Using the smaller size allows us to train over $30$ models and, thus, more accurately estimate the impact of various design choices.
Training protocol
Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer, 2023). Subsequently, we continue training using $1$B tokens with the context extended to $32$K on a mixture of the original RedPajama data (TogetherComputer, 2023) and long context data created using SPLiCe/Baseline. The amount of pre-training tokens is based on scaling laws from (Hoffmann et al., 2022) and constants calculated for GPT-like models in (Karpathy, 2022).
We employ the Focused Transformer (FoT) objective Tworkowski et al. (2023) for context extension (unless stated otherwise). This approach is motivated by practical factors, viz. training with short context length expedites the process, while context scaling can be achieved by finetuning on a relatively small amount of tokens, as demonstrated by (Chen et al., 2023; Tworkowski et al., 2023). Loosely inspired by (Ouyang et al., 2022; Rozière et al., 2023), in the latter phase, long context data (i.e., prepared with SPLiCe) constitutes half of the mixture. We also check how SPLiCe works with other context extension methods in Appendix B.2, and 64K context length in Appendix J.
Evaluation
We measure perplexity on held-out portions of the arXiv (Azerbayev et al., 2022) and StarCoder (Li et al., 2023b) datasets employing a context length of $32$K. The selection of these datasets is motivated by the fact that they can benefit from long-context information as demonstrated in (Chen et al., 2023; Li et al., 2023b). For example, functions are often re-utilized, and similar terms are employed across papers. We exclude documents with fewer than $32$K tokens and truncate those exceeding this length. In Appendix B.1 we evaluate our models using short context data confirming no performance degradation with respect to the base model. For information regarding the training and evaluation data, see Appendix G.
[…]
@@
C SPLiCe BM25 5.46 (.09) 3.20 (.17) 2.81 (.12) 2.22 (.11) 2.94 (.13) 3.10 (.13)
SPLiCe Cont 5.48 (.07) 3.22 (.15) 2.82 (.11) 2.23 (.10) 2.96 (.11) 3.11 (.12)
SPLiCe Repo 5.47 (.08) 3.22 (.15) 2.83 (.10) 2.24 (.09) 2.96 (.11) 3.12 (.11)
Baseline 5.55 3.37 2.93 2.33 3.07 3.23
@@
Baseline 5.550 $\pm$ .002 2.931 $\pm$ .008 3.073 $\pm$ .006 3.228 $\pm$ .005
3.4 Summary of the results
Structuring data improves performance. We observe (see Tables 15) that all tested variants of SPLiCe outperform Baseline, often significantly, regardless of the training and eval dataset. This indicates that the structure of the data is important for long-context training.
The positive effects transfer between domains. We observe, see Tables 2 and 5, that the positive effects of SPLiCe transfer between domains; when training on code and evaluating on Qasper/arXiv and when training on web data and evaluating on code. This is a very interesting effect, which indicates that training induces generic long-context capabilities. Following that, we conjecture that better data mixtures and synthetic long-context data could bring further benefits.
SPLiCe improves the performance of various context extension methods SPLiCe only changes the input structure, and conjecturally, it should work with any context extension method. Our main experiments use the Focused Transformer (FoT) approach (Tworkowski et al., 2023). FoT extends the context only in a couple of attention layers and does not change the positional embeddings of the remaining layers. In Table 9 in Appendix B.2, we confirm that SPLiCe integrates seamlessly with methods that extend the context in all layers. Specifically, we test adjusting Rotary Positional Encodings according to either the recipe as in CodeLlama (Rozière et al., 2023), or YaRN (Peng et al., 2023), or left without changes.
Training is stable. We prepared three subsets of the C code dataset and ran the tested methods on each of them. In Table 5, we report the mean perplexity and its standard deviation. We observe that the differences between the subsets are minimal, which indicates training stability and confirms the statistical significance of our results.
@@ 4 Related work
There is an increasing number of works aiming to study the role of data in LLM training in detail. For instance, (Levine et al., 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training examples leads to better sentence embeddings and improves open-domain question-answering performance. Another work (Gu et al., 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al., 2022). Work (Chan et al., 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities. Similarly, Han et al. (2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
Our methodology diverges from these works in several key ways. Firstly, we focus on the document-level context during the training phase and long context performance, as opposed to sentence-level (Levine et al., 2022) or paragraph-level (Gu et al., 2023) granularity. We demonstrate the efficacy of our approach in large-scale language modeling, specifically with OpenLLaMA $3$B, $7$B, and CodeLlama $13$B. Secondly, we construct a tree structure of related documents through BM25/Contriever-MSMARCO retrieval and linearize this structure to form long-context examples. This allows for greater control over example coherence compared to relying solely on natural data structures like repository-level code. The gradient-based method presented in Han et al. (2023) can be broadly associated with retrieval used in our work; however, it differs in scale and granularity.
Concurrently to our research, (Shi et al., 2023b) showed a method for preparing the training data that closely resembles SPLiCe with default settings ($k=1$ and the identity as Order). The key differences are that (Shi et al., 2023b) focuses on training from scratch while we aim to achieve long context capabilities with short and cheap fine-tuning. Next, we use longer context lengths, up to $32$K tokens, whereas (Shi et al., 2023b) utilizes $8$K. Finally, we provide detailed analysis regarding the design choices, such as the benefits of data reordering in Section 3.2.4 and different values of $k$ in Appendix C.1.
@@ 5 Limitations and future work
We show that structuring the training data is a viable way of improving the model’s performance. The presented method, SPLiCe, can viewed as a general framework for organizing the documents into training examples. This opens multiple further research avenues.
Degree of relevance In this work, SPLiCe constructed training examples by concatenating most matching documents (according to BM25/Cont/Repo). However, recent results of (Tworkowski et al., 2023) show that introducing unrelated data to the context may help the model learn better representations. We leave the study of how the choice of retriever (in particular, the ratio of related and unrelated documents) affects performance for future work.
In order to prepare the training data, SPLiCe uses each document exactly once (we mask out each used document for further retrieval). However, it is possible that allowing some high-quality documents to occur more than once may be beneficial.
Retrieval granularity Another avenue for future work is to study the granularity of the pieces from which the training examples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training examples from smaller pieces.
Other context extension methods In most cases, our models were trained for a long context using the methodology presented in (Tworkowski et al., 2023). Additionally, we tested three popular context extension methods on a medium scale (Naive, YaRN, and CodeLlama) and tuned a large $7$B model using the CodeLlama approach, preliminary confirming the applicability of SPLiCe. However, we leave more detailed studies with other context extension methods to future work.
Other data sources One of the approaches to training long-context language models is to use conversational data (Li et al., 2023a). This is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, or StackExchange questions and answers) to further increase the number of long-context examples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
Data curation Using highly correlated samples has the potential to result in training instability. However, we noted no performance degradation during our experiments. We leave the study of how SPLiCe integrates with different data types for the future.
@@ 6 Conclusions
In this work, we present SPLiCe, a method of constructing training examples for long-context language models. It utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves performance on downstream tasks and the language modeling abilities of LLMs. We further show that SPLiCe can be used to improve long-context utilization of large-scale models using only short fine-tuning. We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
Acknowledgments and Disclosure of Funding
We gratefully acknowledge that our research was supported with Cloud TPUs from Google’s TPU Research Cloud (TRC). We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Centers: ACK Cyfronet AGH, CI TASK) for providing computer facilities and support within computational grant numbers PLG/2023/016265. Piotr Miłoś was supported by National Science Center Poland under the grant agreement 2019/35/O/ST6/03464.
@@ Appendix A Architecture
The architecture of our models is based on LLaMA [Touvron et al., 2023], and the architectural details can be found in Table 6. Briefly speaking, our architecture is similar to the one introduced in [Vaswani et al., 2017] with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding [Su et al., 2021]. For context extension, we use Focused Transformer (FoT)[Tworkowski et al., 2023], CodeLlama [Rozière et al., 2023] (CL) and YaRN Peng et al. [2023]. Table 7 presents the details about both standard and long-context pretraining. We use AdamW as an optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.95$.
Table 6: Architecture details. Focused Transformer context extension is applied in continued pretraining for $32$K context and evaluation.
@@
Table 7: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B/ parameter OpenLLaMAv2 model [Geng, 2023] and $13$B CodeLlama [Rozière et al., 2023] model. Subscript denotes that parameter was specific for a context extension method with FoT referring to [Tworkowski et al., 2023] and no-FoT to other methods (Naive, YaRN [Peng et al., 2023], CodeLlama [Rozière et al., 2023]).
@@ B.1 Short context evaluation
Table 8: $2$K context perplexity evaluation of models from Section 3.3
@@ B.2 Different context extension methods
We test SPLiCe with different context extension methods. In Table 9, we show that SPLiCe brings improvements also when context is extended in all layers using the naive approach, the CodeLlama [Rozière et al., 2023], and [Peng et al., 2023] RoPe [Su et al., 2021] adjustment method.
[…]
@@ B.3 Detailed results
Tables 10 and 11 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method. Table 10 contains detailed results from training on C# and Python. Table 11 contains results on C averaged across three different subsets of C (for details about the construction of those subsets, see Appendix G). Both tables show that SPLiCe outperforms the Baseline by a significant margin.
@@
Tables 13, 14 consider models trained with YaRN [Peng et al., 2023], CodeLlama [Rozière et al., 2023] and Naive (no adjustment to RoPE) context extension methods. Table 15 shows that a simple artificial extension of example length via random concatenation of documents (C source files) does not help.
@@
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
@@
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
@@
Table 11: To check the statistical significance of our results, we prepare three subsets of C (details in Appendix G and train the models on a 50/50 mixture of RedPajama data (organized in the standard way) and C data organized using one of the methods. Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
Altered Train Data: C
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline SPLiCe Repo
ArXiv 5.463 $\pm$ 0.002 5.477 $\pm$ 0.005 5.550 $\pm$ 0.002 5.474 $\pm$ 0.007
C 2.126 $\pm$ 0.020 2.134 $\pm$ 0.020 2.173 $\pm$ 0.023 2.135 $\pm$ 0.020
C++ 2.396 $\pm$ 0.004 2.403 $\pm$ 0.002 2.467 $\pm$ 0.001 2.403 $\pm$ 0.006
CUDA 2.219 $\pm$ 0.002 2.235 $\pm$ 0.005 2.330 $\pm$ 0.009 2.239 $\pm$ 0.004
C# 1.939 $\pm$ 0.000 1.948 $\pm$ 0.001 2.016 $\pm$ 0.004 1.949 $\pm$ 0.002
Dart 2.141 $\pm$ 0.002 2.155 $\pm$ 0.004 2.268 $\pm$ 0.009 2.156 $\pm$ 0.002
Erlang 2.665 $\pm$ 0.002 2.680 $\pm$ 0.003 2.769 $\pm$ 0.007 2.676 $\pm$ 0.003
Fortran 3.306 $\pm$ 0.010 3.335 $\pm$ 0.010 3.554 $\pm$ 0.010 $3$$\pm$ 0.012
Go 1.910 $\pm$ 0.001 1.924 $\pm$ 0.007 1.999 $\pm$ 0.002 1.924 $\pm$ 0.006
Groovy 3.009 $\pm$ 0.001 3.026 $\pm$ 0.006 3.147 $\pm$ 0.013 3.025 $\pm$ 0.007
Haskell 3.198 $\pm$ 0.001 3.221 $\pm$ 0.001 3.371 $\pm$ 0.008 3.220 $\pm$ 0.008
Java 2.075 $\pm$ 0.002 2.086 $\pm$ 0.001 2.161 $\pm$ 0.006 2.085 $\pm$ 0.005
Pascal 3.492 $\pm$ 0.026 3.496 $\pm$ 0.019 3.622 $\pm$ 0.021 3.513 $\pm$ 0.014
Python 2.810 $\pm$ 0.002 2.824 $\pm$ 0.001 2.931 $\pm$ 0.008 2.827 $\pm$ 0.006
Mean 3.100 $\pm$ 0.004 3.115 $\pm$ 0.006 3.228 $\pm$ 0.005 3.115 $\pm$ 0.009
@@
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline
@@
Eval/Method SPLiCe BM25 SPLiCe Cont Baseline
@@
Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
@@
Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
@@
Eval/Method SPLiCe BM25 Baseline SPLiCe Repo
@@
Eval/Method SPLiCe BM25 Baseline DomRnd

### v9

@@ 3.4 Detailed Study with Medium Models
In Table 7, Table 7, and Table 8, we present a comprehensive examination of the impact of document packing on long-context performance using $270$M parameter models, showing that SPLiCe brings consistent improvement in long context language modeling. In Table 7, we scale context to 64K and observe even greater benefits over the Example Packing. We also expand our results to 131K and 160K context length in Tables 19 and 20 in Appendix. In Table 7, we show that SPLiCe is quite robust to the non-accurate retriever. What is more results in Table 7 clearly show that SPLiCe is an improvement over Within-Domain Example Packing. In particular, we note that the more noise we add, the closer SPLiCe is to Within-Domain Example Packing (semantically), and that with 100% noise SPLiCe turns into Within-Domain Example Packing (this is because we use SPLiCe to prepare data coming from a single domain).
Training and Evaluation
Initially, we train with the $2$K context length on $6.3$B tokens from RedPajama (TogetherComputer 2023). Subsequently, we fine-tune using $1$B tokens with the context extended to $32$K on a mixture of the original RedPajama data (TogetherComputer 2023) and long context data created using SPLiCe/EP. We employ the Focused Transformer (FoT) (Tworkowski et al. 2023) for context extension (unless stated otherwise). We measure perplexity on held-out portions of the arXiv (Azerbayev, Piotrowski, and Avigad 2022) and StarCoder (Li et al. 2023b) datasets. The selection of these datasets is motivated by the fact that they can benefit from long-context information as demonstrated in (Chen et al. 2023; Li et al. 2023b). We provide additional details in Appendix M.
@@
EP 5.550 $\pm$ .002 2.931 $\pm$ .008 3.073 $\pm$ .006 3.228 $\pm$ .005
Altered Method arXiv Code Code &
Data Haskell Python CUDA All arXiv
SPLiCeCont 4.88 (.13) 2.62 (.17) 2.67 (.15) 2.34 (.17) 2.77 (.17) 2.90 (.17)
SPLiCeRepo 4.88 (.13) 2.62 (.17) 2.68 (.14) 2.35 (.16) 2.77 (.17) 2.90 (.17)
EP 5.01 2.79 2.82 2.51 2.94 3.07
[…]
Method SPLiCe EP
Eval Data/Noise 0% 10% 25% 50% 75% 90% -
arXiv 5.46 5.47 5.48 5.50 5.53 5.55 5.55
Code Python 2.81 2.82 2.83 2.86 2.89 2.92 2.93
All 2.94 2.95 2.97 3.01 3.04 3.06 3.07
Code & arXiv 3.10 3.11 3.13 3.16 3.19 3.22 3.23
Table 7: We test the robustness of SPLiCe to noisy retriever. We achieve this by preparing data using BM25 retriever that with probability $p$ returns a random document instead of the most related one. We note that SPLiCe is quite robust and only with $p=0.9$ approaches Example Packing.
Altered Method arXiv Code Code
Data Python All arXiv
25% SPLiCe 5.43 2.91 3.08 3.22
25% EP 5.49 3.00 3.19 3.34
50% SPLiCe 5.46 2.81 2.94 3.10
50% EP 5.55 2.93 3.07 3.23
75% SPLiCe 5.61 2.80 2.86 3.05
75% EP 5.73 2.94 3.04 3.23
Table 8: We note that SPLiCe beats the EP perplexity when trained with various proportions of SPLiCeBM25/EP prepared C data (the remaining data is unaltered RedPajama).
3.5 Properties of SPLiCe Generated Data
SPLiCe conjecturally falls into the framework presented in (Chan et al. 2022), which shows that the distributional properties of the training data affect the in-context capabilities of transformer models. In particular, it indicates the importance of ”burstiness”, i.e., a flatter frequency distribution with a relatively higher mass on the rare, long-tail tokens appearing in a sequence. In Table 9, we show that SPLiCe increases the burstiness of the training data (measured in terms of Zipf’s coefficient of token frequency) in comparison to the Example Packing.
Training Data Method Zipf’s Coefficient
C SPLiCeBM25 ${1.512}_{(0.055)}$
EP ${1.593}_{(0.025)}$
StackEx SPLiCeBM25 ${1.643}_{(0.026)}$
EP ${1.664}_{(0.013)}$
Table 9: Zipf’s coefficient of token frequency on EP and SPLiCe along with standard deviation. A lower Zipf’s coefficient represents a more significant burstiness property.
@@ 4 Related Work
There is an increasing number of works aiming to study the role of data in LLM training in detail. For instance, (Levine et al. 2022) developed a theory and demonstrated empirically that incorporating non-adjacent but semantically related sentences in training samples leads to better sentence embeddings and improves open-domain question-answering performance. Another study by (Gu et al. 2023) introduced a pretraining framework grounded on the idea that text documents often include intrinsic tasks. They showed that this approach substantially boosts in-context learning. Additionally, there is existing work on training long-context language models using repository-level code data, such as (Wu et al. 2022). Work of (Chan et al. 2022) identifies the training data’s distributional properties that affect transformer models’ in-context capabilities. Similarly, (Han et al. 2023) constructs small-scale data using an iterative gradient approach and shows that such data improve in-context performance.
Our methodology diverges from these works in several key ways. First, while prior studies have focused on sentence-level (Levine et al. 2022) or paragraph-level (Gu et al. 2023) granularity, we emphasize document-level context during training, specifically targeting long-context performance. We validate our approach in large-scale language modeling, using models such as OpenLLaMA $3$B, $7$B, and CodeLlama $13$B. Second, we construct a tree structure of related documents using BM25/Contriever-MSMARCO retrieval, which we then linearize to form long-context samples. This approach allows for greater control over the coherence of samples, compared to relying solely on natural data structures like repository-level code. While the gradient-based method in (Han et al. 2023) shares similarities with our retrieval-based approach, our method scales to larger datasets and operates at a different granularity.
Concurrently with our research, (Shi et al. 2024) introduced a method for preparing training data that shares similarities with SPLiCe, particularly in its default settings ($k=1$ with identity as Order). However, while their approach focuses on training models from scratch, our work demonstrates that long-context capabilities can be effectively achieved through short and cost-efficient fine-tuning. In addition to this distinction, we employ significantly longer context lengths, extending above 64K tokens compared to the 8K tokens used in (Shi et al. 2024), which allows for more comprehensive context handling. Furthermore, we provide an in-depth analysis of our design choices, such as the advantages of data reordering (detailed in Appendix J) and the impact of varying $k$ values (Appendix C.1). These analyses underline the effectiveness and flexibility of our approach. Our findings are especially pertinent in the context of recent research on the ”Physics of Language Models” (Allen-Zhu and Li 2024), which discusses the limitations of fine-tuning. Despite these limitations, we show that SPLiCe offers substantial and quantifiable improvements even with relatively brief fine-tuning, providing a practical advantage in enhancing long-context capabilities.
@@ 5 Limitations and Future Work
We show that structuring the training data is a viable way of improving the model’s long-context performance. The presented method, SPLiCe, can be viewed as a general framework for organizing the documents into training samples. This opens multiple further research avenues.
Retrieval Granularity Another avenue for future work is to study the granularity of the pieces from which the training samples are constructed. In this work, we focus on the document-level granularity. However, it is possible to construct training samples from smaller pieces.
Other Data Sources One of the approaches to training long-context language models is to use conversational data (Li et al. 2023a), which is complementary to our method. SPLiCe can utilize data that already exists in vast quantities and can be easily applied to different types of text (like code, Wikipedia articles, or StackExchange) to further increase the number of long-context samples. We leave researching how SPLiCe integrates with other methods for preparing the long-context data as future work.
Data Curation Using highly correlated samples has the potential to result in training instability. However, we noted no performance degradation during our experiments. We leave the study of how SPLiCe integrates with different data types for the future. In particular, in our studies, the datasets used were reasonably deduplicated.
Neural Retriever In our work, we have utilized Contriever (Izacard et al. 2022) in a zero-shot setup, using the first $512$ tokens to generate the embedding. On the other hand, BM25 had access to all the document content. Further study is required to determine whether SPLiCe can additionally significantly benefit from properly tuned neural retrievers. In particular, in our case, Contriever tended to produce samples consisting of fewer repositories than BM25. We leave this for future work.
@@ 6 Conclusions
In this work, we present SPLiCe, a method of constructing training samples for long-context language models. It utilizes BM25/Contriever-MSMARCO to find relevant documents and feed them to the model in a structured manner. We show that SPLiCe improves performance on downstream tasks and the language modeling abilities of LLMs. We further show that SPLiCe can be used to improve long-context utilization of large-scale models using only short fine-tuning. We believe that our work indicates multiple interesting research directions for improving the performance of long-context language models with structured data.
Ethical Statement
Our work develops a generic technique that allows for improving context utilization in language models via low-cost fine-tuning. However, we note that it does not create any new threats, but only exacerbates existing ones. Therefore, we refer to the existing literature on the broader impact of language models, such as (Borgeaud et al. 2022).
Acknowledgments
We are thankful for the TPU Research Cloud program, which was instrumental to our research by providing significant computational resources. Parts of the project were realized using the resources of IDEAS NCBR.
@@ Appendix A Architecture
The architecture of our models is based on LLaMA (Touvron et al. 2023), and the architectural details can be found in Table 10. Briefly speaking, our architecture is similar to the one introduced in (Vaswani et al. 2017) with a few standard changes. First, we use only the decoder without the encoder part. Secondly, we perform RMSNorm before the input of both the attention and feed-forward modules. Thirdly, we use the LLaMA FeedForward module. Additionally, we use Rotary Position Embedding (Su et al. 2021). For context extension, we use Focused Transformer (FoT)(Tworkowski et al. 2023), CodeLlama (Rozière et al. 2023) (CL) and YaRN (Peng et al. 2023). Table 11 presents the details about both standard and long-context pretraining/fine-tuning. We use AdamW as an optimizer, with $\beta_{1}=0.9$ and $\beta_{2}=0.95$.
@@
Table 10: Architecture details. Focused Transformer context extension is applied in fine-tuning for $32$K context and evaluation.
@@
Table 11: Training details. We pretrain a custom $270$M parameter model and take a pretrained $3$B/$7$B/ parameter OpenLLaMAv2 model (Geng 2023) and $13$B CodeLlama (Rozière et al. 2023) model. Subscript denotes that parameter was specific for a context extension method with FoT referring to (Tworkowski et al. 2023) and no-FoT to other methods (Naive, YaRN (Peng et al. 2023), CodeLlama (Rozière et al. 2023)).
@@ B.1 Short Context Evaluation
Table 12: $2$K context perplexity evaluation of models from Section 3.4
@@ B.2 Different Context Extension Methods
We test SPLiCe with different context extension methods. In Table 13, we show that SPLiCe brings improvements also when context is extended in all layers using the naive approach, the CodeLlama (Rozière et al. 2023), and (Peng et al. 2023) RoPe (Su et al. 2021) adjustment method.
@@ B.3 Detailed Results
Tables 15 and 14 show the results of training $270$M parameter model for $32$K context on a 50/50 mixture of RedPajama data (organized in a standard way) and code data organized using a specified method. Table 15 contains detailed results from training on C# and Python. Table 14 contains results on C averaged across three different subsets of C (for details about the construction of those subsets, see Appendix F). Both tables show that SPLiCe outperforms the EP by a significant margin.
@@
Tables 17, 19 consider models trained with YaRN (Peng et al. 2023), CodeLlama (Rozière et al. 2023) and Naive (no adjustment to RoPE) context extension methods. Table 21 shows that a simple artificial extension of example length via random concatenation of documents within domain (C source files) does not help.
Altered Train Data: C
Eval/Method SPLiCeBM25 SPLiCeCont EP SPLiCeRepo
ArXiv 5.463 $\pm$ 0.002 5.477 $\pm$ 0.005 5.550 $\pm$ 0.002 5.474 $\pm$ 0.007
C 2.126 $\pm$ 0.020 2.134 $\pm$ 0.020 2.173 $\pm$ 0.023 2.135 $\pm$ 0.020
C++ 2.396 $\pm$ 0.004 2.403 $\pm$ 0.002 2.467 $\pm$ 0.001 2.403 $\pm$ 0.006
CUDA 2.219 $\pm$ 0.002 2.235 $\pm$ 0.005 2.330 $\pm$ 0.009 2.239 $\pm$ 0.004
C# 1.939 $\pm$ 0.000 1.948 $\pm$ 0.001 2.016 $\pm$ 0.004 1.949 $\pm$ 0.002
Dart 2.141 $\pm$ 0.002 2.155 $\pm$ 0.004 2.268 $\pm$ 0.009 2.156 $\pm$ 0.002
Erlang 2.665 $\pm$ 0.002 2.680 $\pm$ 0.003 2.769 $\pm$ 0.007 2.676 $\pm$ 0.003
Fortran 3.306 $\pm$ 0.010 3.335 $\pm$ 0.010 3.554 $\pm$ 0.010 $3$$\pm$ 0.012
Go 1.910 $\pm$ 0.001 1.924 $\pm$ 0.007 1.999 $\pm$ 0.002 1.924 $\pm$ 0.006
Groovy 3.009 $\pm$ 0.001 3.026 $\pm$ 0.006 3.147 $\pm$ 0.013 3.025 $\pm$ 0.007
Haskell 3.198 $\pm$ 0.001 3.221 $\pm$ 0.001 3.371 $\pm$ 0.008 3.220 $\pm$ 0.008
Java 2.075 $\pm$ 0.002 2.086 $\pm$ 0.001 2.161 $\pm$ 0.006 2.085 $\pm$ 0.005
Pascal 3.492 $\pm$ 0.026 3.496 $\pm$ 0.019 3.622 $\pm$ 0.021 3.513 $\pm$ 0.014
Python 2.810 $\pm$ 0.002 2.824 $\pm$ 0.001 2.931 $\pm$ 0.008 2.827 $\pm$ 0.006
Mean 3.100 $\pm$ 0.004 3.115 $\pm$ 0.006 3.228 $\pm$ 0.005 3.115 $\pm$ 0.009
Table 14: To check the statistical significance of our results, we prepare three subsets of C (details in Appendix F) and train the models on a 50/50 mixture of RedPajama data (organized in the standard way) and C data organized using one of the methods. Note that the standard deviation is much lower than the perplexity improvements from using SPLiCe.
@@
Eval/Method SPLiCeBM25 SPLiCeCont EP SPLiCeRepo
@@
Eval/Method SPLiCeBM25 SPLiCeCont EP SPLiCeRepo
@@
Eval/Method SPLiCeBM25 SPLiCeCont EP
@@
Eval/Method SPLiCeBM25 SPLiCeCont EP
@@
Eval/Method SPLiCeBM25 EP SPLiCeRepo
@@
Eval/Method SPLiCeBM25 EP SPLiCeRepo
@@
Eval/Method SPLiCeBM25 EP SPLiCeRepo
@@
SPLiCeBM25 SPLiCeCont Baseline
Eval/Method
ArXiv 4.55 4.56 4.71
C 2.02 2.03 2.12
C++ 1.99 1.99 2.13
CUDA 1.97 1.98 2.14
C# 1.58 1.58 1.66
Common Lisp 2.83 2.84 3.08
Dart 2.13 2.14 2.28
Emacs Lisp 5.90 5.92 6.26
Erlang 1.73 1.74 1.83
Fortran 2.75 2.76 3.06
Go 1.76 1.76 1.87
Groovy 2.32 2.32 2.41
Haskell 2.13 2.14 2.28
Java 1.68 1.69 1.81
Pascal 3.82 3.81 4.05
Python 2.49 2.51 2.63
Mean 2.60 2.61 2.77
Table 19: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C# code with longer $131$K context. We note that SPLiCe still brings significant benefits here, in particular when compared with $32$K and $64$K setups.
SPLiCeBM25 SPLiCeCont Baseline
Eval/Method
ArXiv 4.36 4.36 4.51
Python 2.73 2.74 2.85
Mean 3.54 3.55 3.68
Table 20: Perplexity for training on a $/50$ 50 data mixture of RedPajama and C# code with longer $131$K context and evaluating on $160$K context. We note that FoT uses positional encodings that allow for such extrapolation.
@@
Eval/Method SPLiCeBM25 EP WithinDomEP
@@
SPLiCeBM25 SPLiCeCont Baseline Repo
Eval/Method
ArXiv 4.86 4.88 5.01 4.88
C 2.17 2.19 2.28 2.19
C++ 2.29 2.30 2.43 2.30
CUDA 2.32 2.34 2.51 2.35
C# 1.72 1.73 1.81 1.73
Common Lisp 2.90 2.95 3.15 2.96
Dart 2.07 2.09 2.22 2.09
Emacs Lisp 7.40 7.43 7.84 7.42
Erlang 2.10 2.11 2.23 2.11
Fortran 3.15 3.19 3.50 3.20
Go 1.82 1.83 1.92 1.83
Groovy 2.54 2.59 2.76 2.58
Haskell 2.60 2.62 2.79 2.62
Java 1.83 1.85 1.96 1.84
Pascal 3.62 3.64 3.82 3.62
Python 2.66 2.67 2.82 2.68
Mean 2.88 2.90 3.07 2.90
[…]